OpenStudy (anonymous):

help?! Write the equation in slope-intercept form. What are the slope and y-intercept? -2x-11y=5

OpenStudy (anonymous):

basically solve for y to start

OpenStudy (anonymous):

You want to get it to slope-intercept form which is y=mx+b

OpenStudy (anonymous):

would the answer be y = 2/11x –5/11 ; slope:2/11 ; y-intercept: –5/11

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or would the answer be y = -2/11x –5/11 ; slope: -2/11 ; y-intercept: –5/11

OpenStudy (anonymous):

no slope is +2/11x

OpenStudy (anonymous):

you get division by 2 negatives which is a positive

OpenStudy (anonymous):

ooh okay! so the first answer i posted was correct?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thank you for your help!

OpenStudy (anonymous):

Sure thing